\begin{thebibliography}{00}


\bibitem{BKLS}
S. Bouarroudj, A. Krutov, D. Leites, I. Shchepochkina,
``Non-degenerate invariant (super)symmetric bilinear forms on simple Lie
(super)algebras", Algebras and Repr. Theory.  21(5) (2019), 897--941;
% \url{https://doi.org/10.1007/s10468-018-9802-8};
%\texttt
{arXiv:1806.05505.}


\bibitem{BG} K.A. Brown, I. Gordon,
``Poisson orders, symplectic reflection algebras and representation theory'',
 J.~Reine Angew. Math. 559 (2003), 193 -- 216;
arXiv:math/0201042v2 [math.RT].

\bibitem{sra} P. Etingof and V. Ginzburg,
              ``Symplectic reflection algebras, Calogero--Moser space, and
              deformed Harish--Chandra homomorphism'',
              Inv. Math. 147 (2002), 243 -- 348.

\bibitem{KT} S.E.~Konstein and I.V.~Tyutin,
             ``Traces on the Superalgebra of Observables of Rational Calogero Model
             based on the Root System'',
             Journal of Nonlinear Mathematical Physics, 20:2 (2013), 271 -- 294;
             arXiv:1211.6600; arXiv:math-ph/9904032.

\bibitem{KT7} S.E. Konstein and I.V. Tyutin, ``Ideals generated by traces or by supertraces
in the symplectic reflection algebra $H_{1,\nu} (I_2(2m + 1))$",
Journal of Nonlinear Mathematical Physics, 24:3  (2017), 405-425,
DOI: 10.1080/14029251.2017.1341702;
arXiv:1612.00536. 

\bibitem{KT20} S.E. Konstein and I.V. Tyutin,  ``Connection between the ideals generated
by traces and by supertraces in the superalgebras of observables of Calogero models",
Journal of Nonlinear Mathematical Physics, 27:1 (2020), 7-11, DOI: 10.1080/14029251.2020.1684005;
arXiv:math-ph/1909.02781.

\bibitem{stek} S.E. Konstein and R. Stekolshchik,
              \lq\lq Klein operator and the Number of Traces and Supertraces
              on the Superalgebra of Observables of Rational Calogero Model
              based on the Root System\rq\rq,
              Journal of Nonlinear Mathematical Physics,
              Vol. 20:2 (2013), 295 -- 308.


\bibitem{KLLS}
A. Krutov, A. Lebedev, D. Leites, I. Shchepochkina,
``Non-degenerate invariant symmetric bilinear forms on simple Lie superalgebras
in characteristic 2",
Oberwolfach preprint OWP 2020-02,\\
http://publications.mfo.de/handle/mfo/3697.


\bibitem{IL} I. Losev, \lq\lq Completions of symplectic reflection algebras\rq\rq,
             Selecta Math., 18 (2012), N1, 179-251; arXiv:1001.0239v4.

\bibitem{pass} D.S. Passman, {\it Infinite Crossed Products},
               Pure and Applied Math vol. 135, Academic Press, San Diego, 1989.




\end{thebibliography}

\documentclass[12pt]{article}

\usepackage{euscript,amsmath, amssymb, amsfonts}
\usepackage{color}

\newcommand{\HH}{\mathcal H }
\newcommand{\HG}{$\mathcal H$}

\newcommand{\G}{{\mathbb C}[I_2(2m+1)]}
\newcommand{\spur}{sp^{\varkappa} }

\newcommand{\F}{F_0^{sp^{\varkappa}} }

\newcommand{\m}{\mathfrak {s} }

\newcommand{\Lo}{\mu L_0 }

\newcommand{\be}{\begin{equation}}
\newcommand{\ee}{\end{equation}}
\newcommand{\bee}{\begin{eqnarray}}
\newcommand{\eee}{\end{eqnarray}}
\newcommand\nn{\nonumber \\}

\newcommand\defeq{:=}

\newcounter{theorem}
\newcommand{\theorem}{\par\refstepcounter{theorem}
           {\bf Theorem \arabic{section}.%
           \arabic{theorem}. }}
\renewcommand\thetheorem{\thesection.%
   \arabic{theorem}}
\makeatletter \@addtoreset{theorem}{section}

\newcounter{corollary}
\newcommand{\corollary}{\par\refstepcounter{theorem}
           {\bf Corollary \arabic{section}.%
           \arabic{theorem}. }}

\makeatletter \@addtoreset{corollary}{section}

\newcounter{lemma}

\makeatletter \@addtoreset{lemma}{section}

\newcounter{proposition}
\newcommand{\proposition}{\par\refstepcounter{theorem}
           {\bf Proposition \arabic{section}.%
           \arabic{theorem}. }}
\renewcommand\theproposition{\thesection.%
  \arabic{theorem}}
\makeatletter \@addtoreset{proposition}{section}

\newcounter{conjecture}
\newcommand{\conjecture}{\par\refstepcounter{theorem}
           {\bf Conjecture \arabic{section}.%
           \arabic{theorem}. }}

\makeatletter \@addtoreset{conjecture}{section}

\newcounter{remark}
\newcommand{\remark}{\par\refstepcounter{theorem}
           {\bf Remark \arabic{section}.%
           \arabic{theorem}. }}
\renewcommand\theremark{\thesection.%
     \arabic{theorem}}
\makeatletter \@addtoreset{remark}{section}

\newcounter{definition}
\newcommand{\definition}{\par\refstepcounter{theorem}
           {\bf Definition \arabic{section}.%
           \arabic{theorem}. }}
\renewcommand\thedefinition{\thesection.%
    \arabic{theorem}}
\makeatletter \@addtoreset{definition}{section}

\newenvironment{proof}[1][Proof]{\noindent\textsf{#1.\ }}
{\hfill {\small $\square$}}

\makeatletter \@addtoreset{equation}{section}
\def\theequation{\thesection.\arabic{equation}}

\begin{document}


\sloppy \title
 {
Ideals generated by traces or by supertraces in the symplectic
reflection algebra $H_{1,\nu}(I_2(2m+1))$ II
}

\author
 {
  I.A. Batalin%
\thanks{ I.E. Tamm Department of Theoretical Physics,
          P.N. Lebedev Physical Institute, RAS
          119991, Leninsky prosp., 53, Moscow, Russia.}
\thanks{E-mail: batalin@lpi.ru}
 ,
 S.E. Konstein$^*$%
\thanks{E-mail: konstein@lpi.ru  (corresponding author) }
 ,
 I.V. Tyutin$^*$%
\thanks{E-mail: tyutin@lpi.ru}
}

\date{
}

\maketitle
\thispagestyle{empty}

\begin{abstract}

The algebra $\mathcal H:= H_{1,\nu}(I_2(2m+1))$ of observables of the Calogero model based on the root system
$I_2(2m+1)$ has an $m$-dimensional space of traces and an $(m+1)$-dimensional space of supertraces.
In the preceding paper we found
all values of the parameter $\nu$
 for which either the space of traces contains a~degenerate nonzero trace $tr_{\nu}$ or
the space of supertraces contains a~degenerate nonzero supertrace $str_{\nu}$
and, as a~consequence, the algebra $\mathcal H$ has
two-sided ideals: one consisting of all vectors
in the kernel of the form
$B_{tr_{\nu}}(x,y)=tr_{\nu}(xy)$ or another consisting of all vectors in the kernel of
the form
$B_{str_{\nu}}(x,y)=str_{\nu}(xy)$.
We noticed that if $\nu=\frac z {2m+1}$, where $z\in \mathbb Z \setminus (2m+1) \mathbb Z$, then there exist both a
degenerate trace and a~degenerate supertrace on $\mathcal H$.

Here we prove that the ideals determined by these degenerate
forms coincide.

\end{abstract}

{\footnotesize {\bf Keywords:}
   symplectic reflection algebra,
   trace,
   supertrace,
   ideal,
   dihedral group}



\section{Introduction}
This paper is a continuation of \cite{KT7}; we advise the reader to recall \cite{KT7}.

\subsection{Definitions}

Let $\mathcal A$ be an associative $\mathbb Z_2$-graded algebra with unit;
let  $\varepsilon$ denote its parity. All expressions of linear algebra are given
for homogenous elements only and are supposed to be extended to inhomogeneous elements via
linearity.

A linear complex-valued function $tr$ on $\mathcal A$ is called a~\emph{trace}
if $tr(fg-gf)=0$ for all $f,g\in \mathcal A$.
A linear complex-valued function $str$ on $\mathcal A$ is called a~\emph{supertrace}
if ${str(fg-(-1)^{\varepsilon(f)\varepsilon(g)}gf)=0}$ for all $f,g\in \mathcal A$.
These two definitions can be unified as follows.

Let ${\varkappa=\pm 1}$.
A linear complex-valued function  $\spur$ on $\mathcal A$ is called \emph{$\varkappa$-trace}
if ${\spur(fg-\varkappa^{\varepsilon(f)\varepsilon(g)}gf)=0}$ for all $f,g\in \mathcal A$.

Each nonzero  $\varkappa$-trace $\spur$ defines the nonzero
symmetric\footnote{Initially,
we used the term ``(super)symmetric bilinear form"
currently used by many, e.g., in the paper \cite{BKLS},
even in its title. However, in a recent preprint \cite{KLLS},
it is explained that the supersymmetry
$B(v,w)=(-1)^{p(v)p(w)}B(w,v)$ is related with
the isomorphism $V\otimes W\simeq W\otimes V$
of superspaces and has nothing to do with the (anti)symmetry of the bilinear form $B$ on $V=W$.}
bilinear form $B_{\spur}(f,g):=\spur(fg)$.

If $B_{\spur}$ is degenerate, then the set of the vectors of
its kernel is a~proper ideal in $\mathcal A$. We say that the
 $\varkappa$-trace $\spur$ is \emph{degenerate} if the bilinear form
$B_{\spur}$ is degenerate.

\subsection{The goal and structure of the paper}

The simplicity (or, alternatively, existence of ideals) of
Symplectic Reflection Algebras or, briefly, SRA (for definition, see \cite{sra})
was investigated in a~number of papers, see, e.g.,  \cite{BG}, \cite{IL}.
In particular, it is shown that all SRA $H_{1,\nu}(G)$ with $\nu=0$
are simple (see \cite{pass}, \cite{BG}).

It follows from \cite{KT} and \cite{stek} that an associative algebra of observables of
the Calogero model with harmonic term in the potential and with coupling constant $\nu$
based on the root system $I_2(2m+1)$ (this algebra is SRA denoted $H_{1,\nu}(I_2(2m+1))$)
has an $m$-dimensional space of  traces and
an $(m+1)$-dimensional space of supertraces.

We say that the parameter $\nu$ is \emph{singular}, if the algebra $H_{1,\nu}(I_2(n))$ has
a~degenerate trace or a~degenerate supertrace.

In \cite{KT7}, we found
all singular values of $\nu$
for the algebras $\mathcal H := H_{1,\nu}(I_2(n))$ in the case of $n$ odd ($n=2m+1$) and found the
corresponding degenerate traces and supertraces;
the result is formulated in Theorem \ref{th2}.

We
noticed that if $\nu=\frac z {2m+1}$, where $z\in \mathbb Z \setminus (2m+1) \mathbb Z$,
then there exist both a
degenerate trace  and a~degenerate supertrace on $H$.

Denote this degenerate trace by $tr_z$ and the degenerate supertrace by $str_z$

Theorem \ref{th2} proved in \cite{KT7} implies that if $z\in \mathbb Z \setminus n\mathbb Z$, then

(i) the trace given by the formula (\ref{913})  in \cite{KT7} is degenerate and generates the ideal
$\mathcal I_{tr_z}$ consisting of all the vectors
in the kernel of the degenerate form $B_{tr_z}(x,y)=tr_z(xy)$, 

(ii) 
the supertrace (\ref{914}) is degenerate and generates the ideal $\mathcal I_{str_z}$ consisting of all the vectors in the kernel of
the degenerate form ${B_{str_z}(x,y)=str_z(xy)}$.


The goal of this paper is
Theorem \ref{conj9.1}, which proves
{\conjecture\label{con1} (\cite[Conjecture 9.1]{KT7})
$\mathcal I_{tr_z}=\mathcal I_{str_z}$.
}

\vskip 3mm

In Sections \ref{i2}--\ref{spec} we recall the necessary definitions
and
preliminary facts.

\section{The group $I_2(2m+1)$}\label{i2}

Hereafter in this paper, $n=2m+1$.

\definition
The group $I_2(n)$ is a~finite subgroup of the orthogonal group $O(2,\mathbb R)$ generated
by the root system $I_2(n)$.

The group $I_2(n)$ is the symmetry group of a~flat regular $n$-gon;
$I_2(n)$
consists of $n$
reflections $R_k$ and $n$ rotations $S_k$, where $k=0,1,...,2m$. We consider the indices $k$ as
integers modulo $n$.

These elements ($R_k$ and  $S_k$ for all $k$)
satisfy the relations
\be
\nonumber
R_k R_l = S_{k-l},\qquad
S_k S_l = S_{k+l},\qquad
R_k S_l = R_{k-l},\qquad
S_k R_l = R_{k+l}.
\ee
The element $S_0$
is the unit in the group $I_2(n)$.
Obviously, since $n$ is odd,
all the reflections $R_k$
are in the same conjugacy class.

The rotations $S_k$ and $S_l$ constitute a~
conjugacy class if
$k+l=n$.

{
Let
\be
           \nonumber
\lambda:= \exp\left(\frac {2\pi i}{n}\right).
\ee

Let
\be\label{defg}
G:={\mathbb C}[I_2(n)]
\ee
be the group algebra of the group $I_2(n)$.
In $G$, it is convenient to introduce the following basis
\bee
\nonumber
&& L_p:=\frac 1 n \sum_{k=0}^{n-1} \lambda^{kp} R_k,\qquad
Q_p:=\frac 1 n \sum_{k=0}^{n-1} \lambda^{-kp} S_k.
\eee


\section{Symplectic reflection algebra
$H_{1,\nu}(I_2(2m+1))$}\label{srai2}


{\definition
\emph{The symplectic reflection algebra
$\mathcal H :=H_{1,\nu}(I_2(2m+1))$} is the associative algebra
of  polynomials in the noncommuting elements $a^{\alpha}$ and $b^{\alpha}$, where $\alpha=0,1$,
with coefficients in $G$ (see Eq. (\ref{defg})), satisfying the relations%
}
\bee
\nonumber
&L_p a^\alpha = -   b^\alpha L_{p+1} ,\qquad
&L_p b^\alpha = -   a^\alpha L_{p-1} ,\nn
\label{QH}
&Q_p a^\alpha = a^\alpha Q_{p+1} ,\qquad
&Q_p b^\alpha = b^\alpha Q_{p-1} ,\\
&L_k L_l =  \delta_{k+l} Q_l, \qquad &L_k Q_l =  \delta_{k-l} L_l,\nn
&Q_k L_l =  \delta_{k+l} L_l, \qquad & Q_k Q_l =  \delta_{k-l} Q_l,
\nonumber
\eee
\bee
\left[a^\alpha,\,b^\beta\right]&=& \varepsilon^{\alpha\beta} \hfill
\left( 1+ \mu  L_0 \right), \nn
\left[a^\alpha,\,a^\beta\right]&=&\varepsilon^{\alpha\beta}
\mu   L_1 ,\nn
\left[b^\alpha,\,b^\beta\right]&=&\varepsilon^{\alpha\beta}
\mu   L_{-1},
     \nonumber
     \eee
where $\delta_k\defeq \delta_{k0}$, and
$\varepsilon^{\alpha\beta}$ is the skew-symmetric tensor normalized so that $\varepsilon^{0\,1}=1$, and
 \be
 \nonumber
 \text{~~}\mu:=n {\nu}.
\ee

Defining the parity in \HG  \ by setting
\be     \nonumber
\varepsilon(a^\alpha)=\varepsilon(b^\alpha)=1,\ \ \ \varepsilon(R_k)=\varepsilon(S_k)=0,
\ee
we turn this algebra into a~superalgebra.

The algebra $H_{1,\nu}(I_2(2m+1))$ depends on one complex parameter $\nu$.


\section{Subalgebra of singlets}\label{sl2}

Consider the elements%
\footnote
{Here the brackets $\{\cdot ,\cdot\}$ denote anticommutator.
}
$
T^{\alpha\beta}:=\frac 1 2 (\{a^\alpha,\, b^\beta \}+\{b^\alpha,\, a^\beta \})
$
of the algebra \HG,
\ and the inner derivations of  \HG they generate:
\be     \nonumber
D^{\alpha\beta}:\ \ f\mapsto \left[f,T^{\alpha\beta}\right] \ \
\text{for any } f\in\HH.
\ee
It is easy to verify that the linear span of these derivations
is a~ Lie algebra isomorphic to $sl_2$.

\definition
A \emph{singlet} is any element $f\in \HH$ such that $[f,\, T^{\alpha\beta}]=0$ for all $\alpha,\beta$.
The subalgebra
$H^0 \subset \HH$ consisting of all singlets of the algebra
\HG \ is called  the \emph{subalgebra of singlets}.

One can consider the algebra \HG \ as an $sl_2$-module and decompose it into
the direct sum of irreducible submodules.

Observe, that any $\varkappa$-trace is identically zero on all irreducible
$sl_2$-submodules of \HG,  \ except for singlets.

Let the skew-symmetric tensor $\varepsilon_{\alpha\beta}$ be normalized so that $\varepsilon_{0\,1}=1$
and so $\sum_{\alpha} \varepsilon_{\alpha\beta} \varepsilon^{\alpha\gamma}=\delta_{\beta}^{\gamma}$.
We set
\bee     \nonumber
\m:=\frac 1 {4i} \sum_{\alpha,\beta=0,1}
\left(\{a^\alpha,\, b^\beta \}-\{b^\alpha,\, a^\beta \}\right)\varepsilon_{\alpha\beta}.
\eee

{\proposition \label{1.3.7} (\cite[Proposition 4.2]{KT7})
{\it The subalgebra of singlets $H^0$
is the algebra of polynomials in
the element $\m$ with coefficients in the group algebra  $\G$.}
}

The commutation relations of the singlet $\m$ with generators of
the algebra \HG  \ have the form:
\bee
[\m ,\,Q_p] = [\m ,\,S_k] = [T^{\alpha\beta},\,\m]=0,     \nn
\m L_p =-L_p\m, \qquad \m R_k =-R_k\m,          \nonumber  \nn
(\m -i\mu L_0 )a^\alpha = a^\alpha (\m + i+i \mu L_0).       \nonumber
\eee

\section{The form of ideals in \HG\  and in $H^0$ }\label{id}

{\theorem\label{i0} (\cite[Theorem 4.3]{KT7})
Let $\mathcal I$ be a~proper ideal in the algebra \HG, and
$\mathcal I_0 :=\mathcal I \bigcap H^0$.
Then, there exist  nonzero polynomials
$\phi_k^0\in \mathbb C[\m]$, where $k=0,...,n-1$, such that
$\mathcal I_0$ is the span over $\mathbb C[\m]$ of the elements
\bee               \nonumber
\phi_k^0 (\m)Q_k, \qquad \ \phi_{n-k}^0 L_k, \text{ where } k=0,\,...\,,\,n-1
\ \text{and }\ \phi_n^0:=\phi_0^0.
\eee
}

{\proposition\label{ne0} (\cite[Proposition 4.4]{KT7})
If $\mathcal I\subset \HH$ is a~proper ideal,
then
$\mathcal I_0 =\mathcal I \bigcap H^0$ is a~proper ideal in
$H^0$.
}

\definition
For each $p=0, ... , 2m$, we define the ideals $\mathcal J_p$ and $\mathcal J^p$
in the algebra $\mathbb C[\m]$
by setting
\be     \nonumber
\mathcal J_p := \{f\in \mathbb C[\m] \mid \ f(\m)Q_p \in \mathcal I    \},
\qquad \mathcal J^p := \{f\in \mathbb C[\m] \mid \ f(\m)L_p \in \mathcal I    \}.
\ee

{\proposition\label{8.13} (\cite[Proposition 4.7]{KT7}).
 We have
$\mathcal J_p=\mathcal J^{-p}$.
}

{\proposition\label{9.4} (\cite[Proposition 4.8]{KT7}).
 We have $\mathcal J_p\ne 0$ for any $p=0,...,2m$.
}

Since $\mathbb C[\m]$ is a~principal ideal ring, we have the following statement:

 {\corollary \label{col}   For any
$p=0,...,2m$, there exists a~nonzero polynomial $\phi_p^0\in \mathbb C[\m]$
such that $\mathcal J_p =\phi_p^0 \mathbb C[\m]$.
}

Theorem \ref{i0} evidently follows from Corollary \ref{col}.


\section{Generating functions of $\varkappa$-traces}\label{genfuni2}

For each $\varkappa$-trace $\spur$ on \HG, one can define the following
set of generating functions  which allow one to calculate the $\varkappa$-trace
of arbitrary element in $H^0$ via finding the values of the derivatives of these functions with respect to parameter $t$ at zero:
\bee\label{1.33}
&& F_p^{\spur}(t):=\spur(\exp(t(\m-i\Lo))Q_p),
  \\
&& \Psi_p^{\spur}(t):=\spur(\exp(t\m)L_p),   \text{~~where $p=0,...,2m$.}  \nonumber
\eee

Since $L_0 Q_p=0$ for any $p\ne 0$,
it follows from the definition
~(\ref{1.33}) that
\bee\nonumber
F_p^{\spur}(t)&=& \spur(\exp(t \m)Q_p) \text{ if } p\ne 0,
\nn
F_0^{\spur}(t)&=&\spur(\exp(t(\m-i\mu L_0))Q_0).
\nonumber
\eee

It is easy to find $\Psi_p^{\spur}$ for $p\ne 0$.
Since $\m L_q =-L_q\m$ for any $q=0,...,2m$,
we have
\be\label{psiq}
\Psi_q^{\spur}(t)=\spur(\exp(t\m)L_q)=\spur(L_q).
\ee
Next, since $\spur(R_k)$ does not depend on $k$, we have
$\spur(L_p)=0$ for any $p\ne 0$ and
\be\label{psi0}
\Psi_p^{\spur}(t)\equiv 0 \text{ for any } p \ne 0.
\ee

The value of $\spur(L_0)$ will be calculated later, in Section \ref{GAi2-sp}.

We consider also the functions
\[
\Phi_p^{\spur}(t):=\spur(\exp(t(\m+i\Lo))Q_p).
\]
It is easily verified, by expanding the exponential in a~series, that
these functions are related with the functions $F_p^{\spur}$
by the formula
\be     \nonumber
\Phi_p^{\spur}(t)=F_p^{\spur}(t)+2i\Delta_p^{\spur}(t), \\
\text{where } \Delta_p(t)^{\spur}=\delta_p \sin(\mu t)\spur(L_0).
     \nonumber
\ee

The form of generating functions is related with (non)degeneracy of the form
$B_{\spur}$  as described in Proposition \ref{1.7.8} below.


\section{Degeneracy conditions for the $\varkappa$-trace}\label{degcond}

{\proposition\label{1.7.8} (\cite[Proposition 6.1]{KT7}).
The $\varkappa$-trace on the algebra \HG \ is degenerate  if and only if the
generating functions $F_p^{\spur}$ defined by  formula \eqref{1.33}
have the following form
\bee\label{wid1}
F_p^{\spur}(t)=\sum_{j=1}^{j_p} \exp(t \omega_{j,p})\varphi_{j,p}(t),
\eee
where $\omega_{j,p}\in \mathbb C$ and $\varphi_{j,p}\in \mathbb C [t]$ might depend on $\varkappa$.
}


\section{Equations for the generating functions $F_p^{\spur}$}\label{genfun}

In \cite[Eq. (7.1)]{KT7}, the following system of differential equations for the generating functions is obtained:
\bee\label{eqgenfun}
\frac d {dt} F_p^{\spur} - \varkappa e^{it}\frac d {dt}F_{p+1}^{\spur}=iF_p^{\spur} +
\varkappa ie^{it}F_{p+1}^{\spur}+2 \varkappa i\frac d {dt}\left(e^{it}\Delta_{p+1}^{\spur} \right).
\eee

The initial conditions for this system are:
$$
F_p^{\spur}(0) = \spur(Q_p).
$$

To solve the system (\ref{eqgenfun}), we consider its Fourier transform.
Let
\bee     \nonumber
&& \lambda:=e^{2\pi i/(2m+1)}, \\
&& G_k^{\spur}:= \sum_{p=0}^{2m} \lambda^{kp}F_p^{\spur}, \text{ where } k=0,...,2m,
\label{fur}
\\
&& \widetilde{\Delta}_k^{\spur}:=\sum_{p=0}^{2m} \lambda^{kp}\Delta_{p+1}^{\spur}=
\lambda^{-k}\left( \sin(\mu t)\spur(L_0)    \right),
 \text{ where } k=0,...,2m.
     \nonumber
\eee
For the functions $G_k^{\spur}$, we then obtain the equations
\be\label{Geq}
\frac d {dt} G_k^{\spur} = i\frac {\lambda^k+ \varkappa e^{it}}{\lambda^k- \varkappa e^{it}} G_k^{\spur} +
\frac {2i \varkappa \lambda^k }{\lambda^k- \varkappa e^{it}} \frac d {dt}\left(e^{it}\widetilde{\Delta}_k^{\spur}\right)
\ee
with the initial conditions
\be     \label{init}
G_k^{\spur}(0)=\spur(S_k).
\ee

We choose the following form of the solution of the system (\ref{Geq}):
\bee\label{GK}
G_k^{\spur}(t)= \frac {\varkappa e^{it}} {(\varkappa e^{it}-\lambda^k)^2} \lambda^k g_k^{\spur}(t),
\eee
where
\bee\label{1.47g}
\begin{array}{ll}
g_k^{\spur}(t)&=\left(\frac{2}{\mu}(\cos (t\mu) -1)
+{2i}\lambda ^{-k}(\lambda ^{k}-\varkappa e^{it})\sin (t\mu)
\right)\spur(L_0)\\
&+\varkappa\lambda ^{-k}(\varkappa -\lambda ^{k})^{2} \spur(S_k).
\end{array}
\eee

Evidently, this solution satisfies the initial condition (\ref{init}) for each $\varkappa$
and $k$, except for the case where $\varkappa=+1$ and $k=0$.


If $\varkappa = +1$ and $k=0$, then the expression
~(\ref{GK})
for $G_0^{tr}$ has a~removable singularity at $t=0$.
In this case, instead of the condition ${G_0^{tr}(0)=tr(S_0)}$
we consider the condition $\lim_{t \rightarrow 0} G_0^{tr}(t) = tr(S_0)$.

When $\varkappa=+1$ the solution (\ref{GK}) -- (\ref{1.47g}) gives
\be                 \nonumber
G_0^{tr}(t)= \frac { e^{it}} {( e^{it}-1)^2}
\left(\frac{2}{\mu}(\cos (t\mu) -1)
+{2i}(1- e^{it})\sin (t\mu)
\right)tr(L_0),
\ee
and one can easily see that
\be   \nonumber
\lim_{t \rightarrow 0} G_0^{tr}(t) = -\mu tr(L_0).
\ee

It is shown in Subsection \ref{GAi2-tr}
that if $\varkappa=+1$,
then
\be    \nonumber
tr(S_0) = -\mu tr(L_0)
\ee
for any trace $tr$ on \HG.

So, $G_0^{tr}(t)$ satisfies the initial conditions (\ref{init}) also.


In the case where $ \varkappa = -1$, the $\varkappa$-trace is a~supertrace (see \cite{KT}).
In this case, the $m+1$ values $str(S_k)=str(S_{2m+1-k})$ for $k=0,...,m$ completely define
the supertrace on \HG \ (see \cite{stek}).

In the case where $ \varkappa = +1$, the $\varkappa$-trace is a~trace (see \cite{KT}).
In this case, the $m$ values ${tr(S_k)=tr(S_{2m+1-k})}$ for $k=1,...,m$ completely define
the trace on \HG \ (see \cite{stek}).
The value $tr(S_0)$ linearly depends on parameters $tr(S_k)$, where $k=1,...,m$,
and this value is found in Subsection \ref{GAi2-tr} (see Eqs. (\ref{s0}) -- (\ref{9.6})).

\section{Values of the $\varkappa$-trace on $\G$}\label{GAi2-sp}

From \cite{KT7} we have
\be\label{rk}
\spur (R_k) = - \frac {2\mu} {2m+1}
\left(\frac
{1+\varkappa} 2 X^{tr}
+
\frac {1-\varkappa} 2 Y^{str}
\right),
\ee
where
\bee\label{92}
&& X^{tr}:= \sum_{r=1}^{2m}
\sin^2\left(\frac {\pi r}{2m+1}\right) tr (S_r),\\
&& Y^{str}:= \sum_{r=0}^{2m}                            \label{93}
\cos^2\left(\frac {\pi r}{2m+1}\right) str (S_r).
\eee

Below we consider these values for the traces and supertraces separately.

\subsection{Values of the traces ($\varkappa=+1$) on $\G$}\label{GAi2-tr}

The group $I_2(2m+1)$ has $m$ conjugacy
classes without the eigenvalue +1 in the spectrum:
$
\{S_{p},S_{2m+1-p}\}, \text { where } p=1,...,m.
$

By Theorem 2.3 in \cite{KT}, the values of the trace on these conjugacy classes
\be
\nonumber
s_k := tr(S_k),\ \ \text{ where } s_{2m+1-k}=s_k, \ \ k=1,...,m,
\ee
are arbitrary and
completely define the trace on the algebra \HG.
Therefore, the dimension of the space of traces is equal to $m$.

Further,
the group $I_2(2m+1)$ has one conjugacy class with one eigenvalue +1 in its spectrum:
$
\{R_{1},\,...\,,\, R_{2m+1}\}.
$
The value of $tr(R_k)$ is expressed via $s_k$ by formula (\ref{rk}).

Besides, the group $I_2(2m+1)$ has
one conjugacy class with two eigenvalues +1 in its spectrum:~$\{S_0\}$.

The traces on conjugacy classes with two eigenvalues +1 in the spectrum
is calculated in \cite{KT7} using Ground Level Conditions (for their definition, see \cite{KT}):
\bee
\label{s0}
tr(S_{0}) = 2 \nu^{2}(2m+1) X^{tr}.
\eee

We also note that
\be\label{9.6}
tr(L_0)=-\frac {2\mu} {2m+1} X^{tr},\quad
tr (L_p)=0 \text{ for } p\ne 0, \quad
tr(S_0)=-\mu tr(L_0).
\ee


\subsection{Values of the supertraces ($\varkappa=-1$) on $\G$}\label{GAi2-str}

The group $I_2(2m+1)$ has $m+1$ conjugacy
classes without the eigenvalue $-1$ in the spectrum:
\be     \nonumber
\{S_{0}\},\
\{S_{p},S_{2m+1-p}\}, \text { where } p=1,...,m.
\ee
By \cite[Theorem 2.3]{KT}, the values of the supertrace on these conjugacy classes
\be
\nonumber
u_k := str(S_k)=str(S_{2m+1-k}),\ \ \text{ where }  \ \ k=0,...,m,
\ee
are arbitrary parameters that
completely define the supertrace $str$ on the algebra \HG,
and therefore the dimension of the space of supertraces is equal to $m+1$.

Besides,
the group $I_2(2m+1)$ has one conjugacy class with one eigenvalue $-1$ in the spectrum:
$\{R_{1},\,...\,,\, R_{2m+1}\}$.

The supertraces of the conjugacy class with eigenvalue $-1$ in  its spectrum
are given by Eq. (\ref{rk}):
$str(R_{k}) = -2\nu Y^{str}$, where
$k=0,1,...,2m$, and
where $Y^{str}$ is defined by Eq (\ref{93}).


\section{Singular values of the parameter $\mu$}\label{spec}

 The solution Eq (\ref{GK})-(\ref{1.47g}) determines the generation functions of traces and supertraces
on $H^0$ for any trace and any supertrace on \HG\ . Generally speaking, $G_k^{\spur}$ is a~meromorphic
function on $t$, but if $\mu$ and $\spur$ are such that the form $B_{\spur}$ is degenerate,
then $G_k^{\spur}$ is an integer function on $t$ for each $k$. The complete list of such pairs of $\mu$
and $\spur$ is given in Theorem \ref{th2}. For these values of $\mu$ and $\spur$, the functions
$G_k^{\spur}$ are Laurent polynomials in $\exp(it)$.

{\theorem\label{th2} (\cite[Theorem 9.1]{KT7}).
Let $ m\in \mathbb Z$, where $m\geqslant 1$,
and $n=2m+1$.
Then

\textup{1)} The associative algebra $H_{1,\nu}(I_2(n))$
has a~$1$-parameter set of nonzero  traces $tr_z$ such that the symmetric invariant bilinear form
$B_{tr_z}(x,y)=tr_z(xy)$  is degenerate if and only if
$\nu= \frac z n$, where $z\in \mathbb Z \setminus n\mathbb Z$.
These traces are completely defined by their values at
$S_k$ for
$k=1,\dots , m$:
\be\label{913}
tr_z(S_k)= \frac \tau {n \sin^{2}\frac {\pi k}n}\left
(1 - \cos\frac {2\pi k z}{n}\right), \text{ ~~where } \tau\in \mathbb C,\ \tau\ne 0.
\ee
Here $\tau$ is an arbitrary parameter specifying the trace in 1-dimensional space of
 traces.

\textup{2)}  The associative superalgebra $H_{1,\nu}(I_2(n))$
has a~$1$-parameter set of  nonzero supertraces $str_z$ such that the symmetric invariant bilinear form
$B_{str_z}(x,y)=str_z(xy)$  is degenerate if
$\nu= \frac z n$, where $z\in \mathbb Z \setminus n\mathbb Z$.
These supertraces are completely defined by their values at
$S_k$ for
$k=0,\dots , m$:
\bee\label{914}
str_z(S_k)= \frac \tau {n \cos^{2}\frac {\pi k}n}
\left(1 - (-1)^{z}\cos\frac {2\pi k z}{n}\right), \text{ where } \tau\in \mathbb C,\ \tau\ne 0.
\eee
Here $\tau$ is an arbitrary parameter specifying the supertrace in 1-dimensional space of supertraces.

\textup{3)}  The associative superalgebra $H_{1,\nu}(I_2(n))$
has a~$1$-parameter set of nonzero supertraces $str_{1/2}$ such that the symmetric invariant bilinear form
$B_{str_{1/2}}(x,y)=str_{1/2}(xy)$  is degenerate if
$\nu= z + \frac 1 2$, where $z\in \mathbb Z$.
These supertraces are completely defined by their values at
$S_k$ for
$k=0,\dots , m$:
\be             \nonumber
str_{1/2}(S_k)= \frac \tau {n \cos^{2}\frac {\pi k}n}, \text{ where } \tau\in \mathbb C, \ \tau\ne 0.
\ee
Here $\tau$ is an arbitrary parameter specifying the supertrace in 1-dimensional space of supertraces.

\textup{4)}  For all other values of $\nu$, all nonzero traces and supertraces are nondegenerate.
}

\section{Generating functions $F_p^{\spur}$   for the degenerate $\varkappa$-trace}

Let $\mu \in\mathbb Z \setminus n\mathbb Z$. Substitute the solutions
(\ref{913})   for the case $\varkappa=+1$  and
(\ref{914})   for the case $\varkappa=-1$
to Eqs. (\ref{GK})--Eq. (\ref{1.47g}). We obtain
the formula for both values of $\varkappa$
\be\label{g2}
g_k^{\spur}=-\frac{4\tau}{n}\left[
\cos(t\mu)+i\mu\lambda^{-k}(\lambda^k-\varkappa e^{it})\sin(t\mu)
-\varkappa^\mu\cos\frac{2\pi k\mu}{n}
\right].
\ee
Introducing the new variable $y$ instead of $t$
\be\label{y}
y:=\varkappa e^{it}
\ee
we can rewrite Eq. (\ref{g2}) in the form
\be
                       \nonumber
g_k^{\spur}=-\frac{2\tau}{n}\varkappa^\mu
\left[
(y^\mu+y^{-\mu})+\mu\lambda^{-k}(\lambda^k-y)(y^\mu-y^{-\mu})
-2\cos\frac{2\pi k\mu}{n}\right]
\ee
and Eq. (\ref{GK}) in the form
\be\label{GK2}
G_k^{\spur}=\frac{\lambda^k y}{(y-\lambda^k)^2}g_k^{\spur} .
\ee
Now we see that $G_k^{\spur}$ are the Laurent polynomials  in $y$ with the highest degree $\leq |\mu|$
and the lowest degree $\geq 1-|\mu|$.

Note, that the expressions  (\ref{GK2}) are even functions of the parameter $\mu$,
so we can assume that $\mu$ is a~positive integer.

Let $\mu > 0$ in what follows.

Thus, $G_k^{\spur}$ can be expressed in the form
\be\label{GK3}
G_k^{\spur}=\varkappa^\mu \sum_{\ell=\mu}^{1-\mu}\beta_{\ell}^{k}y^{\ell},
\ee
where
the 
$\beta_{\ell}^{k}$
are constants not depending on $\varkappa$ and
not all of them
equal to zero.

Eq (\ref{GK3}) implies that
\be\label{p0}
\beta_{\mu}^k=\frac{2\tau\mu}n.
\ee

Further, Eq. (\ref{fur}) implies
\be
       \nonumber
F_p^{\spur}=\frac 1 n \sum_{k=0}^{2m}\lambda^{-kp}G_k^{\spur}
\ee
and the generating functions $F_p^{\spur}$ have the form
\be\label{Fform}
F_p^{\spur}=\varkappa^\mu \sum_{\ell=\mu}^{-\mu}\alpha_{\ell}^{p}y^{\ell},
\ee
 where the $\alpha_{\ell}^{k}$ are constants not depending on $\varkappa$.
Observe that $F_p^{\spur}$ can be equal to zero for some $p\ne 0$ (e.g., if $\mu=1$, then $F_p^{\spur}=0$ for
each $p\ne 0$),
but $F_0^{\spur}\ne 0$
since Eq (\ref{p0}) implies $\alpha^0_{\mu}=\frac{2\tau\mu}n\ne 0$. Eq (\ref{p0}) implies also that $\alpha^0_{-\mu}=0$.


\section{The generating function $\mathcal F^{\spur}$=$\spur\left(\exp(t\m )Q_0 \right)$
  for the degenerate $\varkappa$-trace}

 Let $\mu \in\mathbb Z \setminus n\mathbb Z$ and $\varkappa$-trace be defined by
Eq (\ref{913}) in the case $\varkappa=+1$  and
by Eq (\ref{914}) in the case $\varkappa=-1$.

In this section we introduce the function
\be
           \nonumber
\mathcal F^{\spur}:=\spur\left(\exp(t \frak s )Q_0 \right)
\ee
and express it via $F_0^{\spur}$.
{\proposition\label{p121}
$\mathcal F^{\spur}$ is an even function of $t$:
\be\label{121}
\mathcal F^{\spur}=\spur(\cosh(t\m)Q_0 .
\ee
}

Indeed, $\mathcal F^{\spur}=\spur(\cosh(t\m)Q_0+\sinh(t\m)Q_0)$ and
$\spur(\sinh(t\m)Q_0)=0$ since
\bee\nonumber
\spur(\sinh(t\m)Q_0)=\spur((\sinh(t\m)L_0) L_0)=\spur(L_0(\sinh(t\m)L_0))=\nn=
\spur((L_0\sinh(t\m))L_0)=\spur((-(\sinh(t\m)L_0))L_0)
=\spur(-\sinh(t\m)Q_0)\nonumber
\eee

Now, decompose $\F$:
\bee
\F&=&\spur \left(e^{t(\m -i\mu L_0)} Q_0\right)=F_{even}+F_{odd}  \text{ where } \nn
F_{even}&=& \spur \left(\sum_{s=0}^{\infty}\frac 1 {(2s)!}(t(\m -i\mu L_0))^{2s} Q_0\right)
=\spur \left(\sum_{s=0}^{\infty}\frac 1 {(2s)!}t^{2s}(\m^2 -\mu^2)^{s} Q_0\right) ,                \label{even}
\\
F_{odd}&=& \spur \left(\sum_{s=0}^{\infty}\frac 1 {(2s+1)!}(t(\m -i\mu L_0))^{2s+1} Q_0\right)=  \nn
       &=&\spur \left(\sum_{s=0}^{\infty}\frac 1 {(2s+1)!}t^{2s+1}(\m^2 -\mu^2)^{s}
       (\m -i\mu L_0)Q_0\right)= \nn
       &=&
        \spur \left(\sum_{s=0}^{\infty}\frac 1 {(2s+1)!}t^{2s+1}(\m^2 -\mu^2)^{s}
       ( -i\mu L_0)Q_0\right) = \nn
&=&
        \sum_{s=0}^{\infty}\frac 1 {(2s+1)!}t^{2s+1}(-\mu^2)^{s}
       ( -i\mu)\spur L_0 =                       \nn
  &=&     \sinh (-i\mu)\spur L_0 =
       -\frac {\varkappa^\mu}{2}(y^{\mu}-y^{-\mu}) \spur L_0     .                 \label{odd}
\eee
Eq. (\ref{Fform}) implies that
\be\label{odd2}
F_{odd}= \frac {\varkappa^\mu} 2
\left(  \sum_{\ell=\mu}^{-\mu}\alpha_{\ell}^{0}y^{\ell} - \sum_{\ell=\mu}^{-\mu}\alpha_{-\ell}^{0}y^{\ell}
\right).
\ee
Comparing Eq. (\ref{odd2})  with Eq. (\ref{odd}) implies
\bee                     \nonumber
\alpha^0_{\ell}&=&\alpha^0_{-\ell}, \qquad \text{ if $\ell \ne \mu$, $\ell\ne -\mu$},  \\
\alpha^0_{\mu}&-&\alpha^0_{-\mu}= -\spur L_0,
\eee
and
\be\label{127}
F_{even}=\frac {\varkappa^{\mu}}{2}\alpha^0_{\mu}(y^{\mu}+y^{-\mu})
+\frac {\varkappa^{\mu}}{2} \sum_{\ell=0}^{\mu-1} \alpha^0_{\ell}(y^{\ell}+y^{-\ell})=
 \alpha^0_{\mu}\cosh(it\mu)
+ {\varkappa^{\mu}} \sum_{\ell=0}^{\mu-1} {\varkappa^{\ell}}\alpha^0_{\ell}\cosh (it\ell) .
\ee

\medskip

{\proposition
\be
\nonumber
\mathcal F^{\spur}(t) =
\alpha^0_{\mu}
+ {\varkappa^{\mu}} \sum_{\ell=0}^{\mu-1} {\varkappa^{\ell}}\alpha^0_{\ell}\cosh \left(t\sqrt{\mu^2-\ell^2}\,\right) .
\ee
}

\medskip

\medskip

\begin{proof}
Taking Proposition \ref {p121}  into account let us decompose Eq (\ref{121}) into the Taylor series:
\be            \nonumber
\mathcal F^{\spur}(t)=\sum_{s=0}^{\infty}a_{2s}\frac {t^{2s}}{(2s)!},
\ee
where $a_{2s}:=\spur(\m^{2s}Q_0)$ for $s=0,1,2,\,...$.

Eq (\ref{even}) implies
\be                 \nonumber
a_{2s}=\left( \frac {d^2}{dt^2} + \mu^2\right)^s F_{even}|_{t=0},
\ee
and Eq (\ref{127}) implies
\be                   \nonumber
a_{2s}=\left\{
\begin{array}{c}
a_{\mu }^{0} +
{\varkappa^{\mu}} \sum_{\ell=0}^{\mu-1} {\varkappa^{\ell}}\alpha^0_{\ell}
\text{ \ \ \ if }s=0 \\
{\varkappa^{\mu}} \sum_{\ell=0}^{\mu-1} {\varkappa^{\ell}}\alpha^0_{\ell}(-\ell^2+\mu^2)^s
 \text{ \ \ \  if }s\ne0    .
\end{array}%
\right.
\ee
So
\be                       \nonumber
\mathcal F^{\spur}(t)=\sum_{s=0}^{\infty}a_{2s}\frac {t^{2s}}{(2s)!}=
\alpha^0_{\mu}
+ {\varkappa^{\mu}} \sum_{\ell=0}^{\mu-1} {\varkappa^{\ell}}\alpha^0_{\ell}\cosh \left(t\sqrt{\mu^2-\ell^2}\,\right)  .
\ee
\end{proof}

\section{Ideals generated by degenerate $\varkappa$-traces}

{Let $\mu=z \in\mathbb Z \setminus n\mathbb Z$ and the $\varkappa$-trace be defined by
Eq (\ref{913}) for the case $\varkappa=+1$  and
by Eq (\ref{914}) for the case $\varkappa=-1$.

These degenerate $\varkappa$-traces are denoted in Theorem \ref{th2} by $tr_z$ and
$str_z$.

Denote the ideals generated by these traces $\spur$ by $\mathcal I^{\varkappa}$;   in $H^0$, consider
the ideals  $\mathcal I_0^{\varkappa} := \mathcal I^{\varkappa}\bigcap H^0$.

Now we can prove Conjecture \ref{con1} (\cite[Conjecture 9.1]{KT7}):

{\theorem\label{conj9.1}
$\mathcal I^{+1}=\mathcal I^{-1}$.
}

To prove Theorem \ref{conj9.1} we use Theorem 4.2 from \cite{KT20} which  in our case implies

{\theorem\label{4.2} (\cite[Theorem 4.2]{KT20})
$\mathcal I^{+1}=\mathcal I^{-1}$ if and only if $\mathcal I^{+1}\bigcap H^0=\mathcal I^{-1}\bigcap H^0$.
}

So, Theorem \ref{conj9.1} follows from 
{\theorem\label{4.3}
$\mathcal I^{+1}\bigcap H^0=\mathcal I^{-1}\bigcap H^0$.
}

\begin{proof}
For degenerate $\spur$, we established the following facts:
\bee             \nonumber
&&F_p^{\spur}(t)=\spur\left(e^{t\m}Q_p \right)=\varkappa^\mu\sum_{\ell=\mu}^{-\mu}
\alpha_{\ell}^{p}\varkappa^{\ell}
e^{it\ell} \text{ \ for \ } p=1,2,...,n-1  ,\\
&&\mathcal F^{\spur}(t)=\spur\left(e^{t\m}Q_0   \right)=
\alpha^0_{\mu}
+ {\varkappa^{\mu}} \sum_{\ell=0}^{\mu-1} {\varkappa^{\ell}}\alpha^0_{\ell}\cosh \left(t\sqrt{\mu^2-\ell^2}\,\right)
\text{\ \ where\ \ } \alpha^0_{\mu}\ne 0,
\nonumber
\eee
and where the $\alpha$-s do not depend on $\varkappa$.


For any $p=1,...,n$, it is easy to find the lowest degree polynomial differential operators with constant
coefficients $D^{\varkappa}_p(d/dt)$
 such that $D^{\varkappa}_p(d/dt) F_p^{\spur}(t)=0$:
\be              \nonumber
D^{\varkappa}_p\left(\frac d{dt}\right)= \begin{cases}
\mathop{\prod}\limits_{\ell=-\mu:\ \alpha_{\ell}^{p}\ne 0}^{\mu}
\left(\frac d {dt} -i\ell \right)&\text{\ \ if $F_p^{\spur}\ne 0$,}\\ 1&
\text{\ \ if $F_p^{\spur} = 0$,}\end{cases}
\ee
and $D^{\varkappa}_0(d/dt)$ such that $D^{\varkappa}_0(d/dt) \mathcal F^{\spur}(t)=0$:
\be                \nonumber
D^{\varkappa}_0\left(\frac d{dt}\right)=\frac d {dt}
\prod_{\ell=0:\ \alpha_{\ell}^{0}\ne 0}^{\mu-1}
\left(\frac {d^2 }{dt^2} -\mu^2+ \ell^2 \right).
\ee
Further, it is a~simple exercise to prove that
\be             \nonumber
D_p^{\varkappa}(\m)Q_p,\ D_p^{\varkappa}(\m)L_{-p} \in \mathcal I^{\varkappa}_0
\text{\ \ \ for any\ \ \ } p=0,\,...\,n-1,
\ee
namely,
\be               \nonumber
B_{\spur}(D_p^{\varkappa}(\m)Q_p,\, f )=B_{\spur}(D_p^{\varkappa}(\m)L_{-p},\, f )=0
\text{\ \ \ for any $f\in H^0$ and $p=0,\,...\,n-1$}.
\ee

Consider, for example, $B_{\spur}(D_0^{\varkappa}(\m)Q_0,\, f)$ for $f=g(\m)Q_p$ and $f=g(\m)L_p$:
\be             \nonumber
\spur(D_0^{\varkappa}(\m)Q_0 g(\m)Q_p)=\spur(D_0^{\varkappa}(\m)Q_0 g(\m)L_p)=0 \text{\ \ for \ \ }p\ne0,
\ee
since $Q_0 Q_p=Q_0 L_p=0$ for $p\ne 0$,
\bee
\spur(D_0^{\varkappa}(\m)Q_0 g(\m)Q_0)=\spur(D_0^{\varkappa}(\m) g(\m)Q_0)&=&
D_0^{\varkappa}\left(\frac d {dt}\right)g\left(\frac d {dt}\right) \spur\left(e^{t\m}Q_0    \right)|_{t=0}=  \nn
&=&g\left(\frac d {dt}\right) D_0^{\varkappa}\left(\frac d {dt}\right) \mathcal F^{\spur}(t)|_{t=0}=0,
\nonumber
\eee
\bee
\spur(D_0^{\varkappa}(\m)Q_0 g(\m)L_0) = \spur(D_0^{\varkappa}(\m) g(\m)L_0)&=&
D_0^{\varkappa}\left(\frac d {dt}\right)g\left(\frac d {dt}\right) \spur\left(e^{t\m} L_0   \right)|_{t=0}=  \nn
&=&g\left(\frac d {dt}\right) D_0^{\varkappa}\left(\frac d {dt}\right) \spur(L_0)=0
\nonumber
\eee
due to Eq (\ref{psiq}) and since the operator $D_0^{\varkappa}\left(\frac d {dt}\right)$
contains the factor $\frac d {dt}$.


Further, it is easy to see that for each of the ideals $\mathcal I_0^{\varkappa}$, where $\varkappa=\pm 1$,
the polynomials $\phi_p^0\in \mathbb C[\m]$ defined in Corollary \ref{col} satisfy the relations
$\phi_p^0(\m)=D_p^{\varkappa}(\m)$ for $p=0,...,n-1$.

So, Theorem \ref{i0} implies that the $\mathbb C [\m]$-span of the $D_p^{\varkappa}(\m)Q_p$ and $D_p^{\varkappa}(\m)L_{-p}$
for $p=0,...,n-1$
is $\mathcal I^{\varkappa}_0$.

Since $D^{+1}_p=D^{-1}_p$, we have $\mathcal I^{+1}_0=\mathcal I^{-1}_0$,
and as result, $\mathcal I^{+1}=\mathcal I^{-1}$.
\end{proof}

\vskip 5mm

\section*{Acknowledgments}
The authors
are grateful to Russian Fund for Basic Research
(grant No.~${\text{20-02-00193}}$)
for partial support of this work.


\setcounter{equation}{0} \def\theequation{A
\arabic{equation}}

\newcounter{appen}
\newcommand{\appen}[1]{\par\refstepcounter{appen}
{\par\bigskip\noindent\large\bf Appendix. 
\medskip }{\bf \large{#1}}}

\renewcommand{\subsection}[1]{\refstepcounter{subsection}
\vskip 3mm{\centerline{\bf 
A\arabic{subsection}. \ #1}}\vskip 3mm}
\renewcommand\thesubsection{A\theappen.\arabic{subsection}}
\makeatletter \@addtoreset{subsection}{appen}

\renewcommand{\subsubsection}{\par\refstepcounter{subsubsection}
{\bf A\arabic{appen}.\arabic{subsection}.\arabic{subsubsection}. }}
\renewcommand\thesubsubsection{A\theappen.\arabic{subsection}.\arabic{subsubsection}}
\makeatletter \@addtoreset{subsubsection}{subsection}


\newcounter{theor}
\renewcommand{\theorem}{\par\refstepcounter{theor}
           {\bf Theorem A.%
           \arabic{theor}. }}
\renewcommand\thetheorem{A.\arabic{theor}}
\makeatletter \@addtoreset{theor}{section}

\renewcommand{\remark}{\par\refstepcounter{theor}
           {\bf Remark A.%
           \arabic{theor}. }}
\renewcommand\theremark{\thesection.%
     \arabic{theor}}
\makeatletter \@addtoreset{remark}{section}

\renewcommand{\definition}{\par\refstepcounter{theor}
           {\bf Definition A.%
           \arabic{theor}. }}
\renewcommand\thedefinition{\thesection.%
    \arabic{theor}}
\makeatletter \@addtoreset{definition}{section}

\renewcommand{\proposition}{\par\refstepcounter{theor}
           {\bf Proposition 
           A.\arabic{theor}. }}
\renewcommand\theproposition{
  A.\arabic{theor}}
\makeatletter \@addtoreset{proposition}{section}

\renewcommand{\theorem}{\par\refstepcounter{theor}
           {\bf Theorem 
           A.\arabic{theor}. }}
\renewcommand\thetheorem{
   A.\arabic{theor}}
\makeatletter \@addtoreset{theorem}{section}




\end{document}